f(x)=cos√sin(tanπx)


1
Expert's answer
2021-05-24T18:50:50-0400

"f(x)=cos\sqrt{sin(tan\pi x)}"

Using chain rule of differentiation

"\dfrac{d[f(x)]}{dx}=-sin\sqrt{sin(tan\pi x)}\times\dfrac{1}{2\sqrt{sin(tan\pi x)}}\times cos(tan\pi x)\times sec(2\pi x)\times \pi"

"f'(x)=-\dfrac{\pi \sin\sqrt{\sin(\tan\pi x)}\space \cos(\tan\pi x)\space \sec(2\pi x) }{2\sqrt{\sin(\tan(\pi x))}}"